OpenStudy (awkwardpanda):

find the value of tan (a-B) if cos a=3/5, sin B=5/13, 0 degrees< a <90degrees, and 0 degrees < B <90 degrees

OpenStudy (awkwardpanda):

@chlobohoe

OpenStudy (anonymous):

tan(a-b) = (tana - tanb)/(1 + tanatanb) a in Q2 : sina > 0 cosa = -3/5 sina = 4/5 tana = -4/3 b in Q2: sinb < 0 sinb = 5/13 cosb = -12/13 tanb = -5/12 tan(a-b) = (-4/3 + 5/12)/(1 + 20/36) tan(a-b) = (-48+15)/(36 + 20) = -33/56 ==Note: the correct solution is -33/56, not +33/56

OpenStudy (anonymous):

btw i love ur name <3 <3 <3

OpenStudy (awkwardpanda):

Thank you! <3